Assuming no change in temperature and pressure, calculate the volume of O2 (in liters) required for the complete combustion of 17.6 L of butane (C4H10):

2C4H10(g) + 13O2(g) → 8CO2(g) + 10H2O(l)

 

1
Expert's answer
2022-03-27T10:31:57-0400

V(C4H10) = 17.6 L;

Vm = 22.4 L/mol;

n(C4H10) = V(C4H10)/Vm = 17.6/22.4 mol;

2C4H10(g) + 13O2(g) → 8CO2(g) + 10H2O(l);

For the chemical reaction:

n(O2) = 13/2 * n(C4H10) = 13/2 * 17.6/22.4;

V(O2) = n(O2) * Vm = 13/2 * 17.6/22.4 * 22.4 = 13/2 * 17.6 = 114.4 L.

Answer: 114.4 L.